# CR-CTC: Consistency regularization on CTC for improved speech recognition

### v2

@@ Abstract
Connectionist Temporal Classification (CTC) is a widely used method for automatic speech recognition (ASR), renowned for its simplicity and computational efficiency. However, it often falls short in recognition performance compared to transducer or systems combining CTC and attention-based encoder-decoder (CTC/AED). In this work, we propose the Consistency-Regularized CTC (CR-CTC), which enforces consistency between two CTC distributions obtained from different augmented views of the input speech mel-spectrogram. We provide in-depth insights into its essential behaviors from three perspectives: 1) it conducts self-distillation between random pairs of sub-models that process different augmented views; 2) it learns contextual representation through masked prediction for positions within time-masked regions, especially when we increase the amount of time masking; 3) it suppresses the extremely peaky CTC distributions, thereby reducing overfitting and improving the generalization ability. Extensive experiments on LibriSpeech, Aishell-1, and GigaSpeech datasets demonstrate the effectiveness of our CR-CTC, which achieves performance comparable to, or even slightly better than, that of transducer and CTC/AED. We release our code at https://github.com/k2-fsa/icefall.
@@
Peaky CTC distributions. CTC models are known for predicting extremely peaky distributions (Graves et al., 2006; Sak et al., 2015), which can be harmful in certain scenarios, such as forced alignment (Huang et al., 2024) and knowledge distillation (Ding et al., 2020). These peaky distributions lead to inaccurate alignments as the model assigns excessive blanks to non-silence frames. To address this, label priors are employed to suppress the peaky distributions, thereby improving the accuracy of forced alignment (Huang et al., 2024). As position mismatches of CTC spikes can hinder knowledge distillation performance, some approaches propose to encourage consistent alignments between the teacher and student (Ding et al., 2020) or to utilize sequence-level distillation (Takashima et al., 2019). Unlike previous work, we demonstrate that peak suppression in CR-CTC can improve the generalization ability of the CTC model.
@@
Consistency regularization. The technique of consistency regularization has demonstrated effectiveness in learning generalizable representations across various learning paradigms, including self-supervised (Chen et al., 2020; Grill et al., 2020; Chen & He, 2021), semi-supervised (Sajjadi et al., 2016; Laine & Aila, 2016; Sohn et al., 2020; Masumura et al., 2020; Chen et al., 2021b; Higuchi et al., 2021; Qi et al., 2022), and supervised (Wu et al., 2021; Gao et al., 2022; Touvron et al., 2023; Heo et al., 2023; Yoon et al., 2024) learning tasks. SimSiam (Chen & He, 2021), for instance, maximizes the similarity between two augmented views of the same image. FixMatch (Sohn et al., 2020) generates pseudo-labels based on high-confidence predictions from weakly augmented images, then trains the model to predict these pseudo-labels using the strongly augmented versions of the same images. Several methods encourage consistency between the predictions of randomly sampled sub-models using drop-based techniques, such as R-Drop (Wu et al., 2021) and cosub (Touvron et al., 2023). The idea of R-Drop has also been applied to CTC models (Graves et al., 2006) for ASR (Gao et al., 2022; Yoon et al., 2024). For example, in (Gao et al., 2022), spatial-temporal dropout is specially designed to construct the sub-models, with consistency regularization enforced exclusively on the CTC spike frames. Similar to our CR-CTC, these methods utilizing consistency regularization are also interpreted as performing self-distillation between the random sub-models (Wu et al., 2021; Touvron et al., 2023).
@@ 3.1 Preliminary: Connectionist Temporal Classification
$\mathcal{L}_{\mathrm{CTC}}(\mathbf{x},\mathbf{y})=-\log\sum_{\bm{\pi}\in\mathcal{B}^{-1}(\mathbf{y})}p(\bm{\pi}|\mathbf{x}).$ (1)
@@ 3.2 Our approach: Consistency-regularized CTC
Figure 1 illustrates the overall architecture of our proposed CR-CTC. It takes as input two different augmented views, $\mathbf{x}^{(a)}$ and $\mathbf{x}^{(b)}$, both derived from the input speech mel-spectrogram $\mathbf{x}$. The two input views are then passed through a shared speech encoder $f$, which estimates the per-frame distributions: $\mathbf{z}^{(a)}=f(\mathbf{x}^{(a)})$, $\mathbf{z}^{(b)}=f(\mathbf{x}^{(b)})$. In addition to computing the CTC losses on both branches: $\mathcal{L}_{\mathrm{CTC}}(\mathbf{x}^{(a)},\mathbf{y})$ and $\mathcal{L}_{\mathrm{CTC}}(\mathbf{x}^{(b)},\mathbf{y})$, we introduce an auxiliary loss (defined in Equation 4) to enforce consistency between $\mathbf{z}^{(a)}$ and $\mathbf{z}^{(b)}$: $\mathcal{L}_{\mathrm{CR}}(\mathbf{z}^{(a)},\mathbf{z}^{(b)})$. The overall loss of the whole model is formulated as:
$\mathcal{L}=\frac{1}{2}(\mathcal{L}_{\mathrm{CTC}}(\mathbf{x}^{(a)},\mathbf{y})+\mathcal{L}_{\mathrm{CTC}}(\mathbf{x}^{(b)},\mathbf{y}))+\alpha\mathcal{L}_{\mathrm{CR}}(\mathbf{z}^{(a)},\mathbf{z}^{(b)}),$ (3)
@@ 4.1 Experimental setup
Implementation details. Our experiments are performed using the icefall framework  [^2] https://github.com/k2-fsa/icefall, with Lhotse toolkit (Żelasko et al., 2021) for data preparation. For regular ASR recipes in icefall, default parameter settings of SpecAugment (Park et al., 2019) include a time warping factor of 80, 2 frequency masking regions with a maximum width of 27, and 10 time masking regions with a maximum width of 100, along with a maximum masking fraction of 15% specifically for time masking. In our CR-CTC systems, we utilize larger amount of time masking through increasing both the number of time masking regions and the maximum masking fraction by a factor of 2.5. Speed perturbation (Ko et al., 2015) with factors 0.9, 1.0 and 1.1 is applied to LibriSpeech (Panayotov et al., 2015) and Aishell-1 (Bu et al., 2017) datasets. The input features are 80-dimensional mel-spectrograms extracted using 25-ms windows with a 10-ms shift. For LibriSpeech and GigaSpeech datasets, we employ 500-class Byte Pair Encoding (BPE) (Sennrich et al., 2016) word pieces as modeling units, while for Aishell-1 dataset, we use 4336-class characters. By default, we set $\alpha$ in Equation 3 to 0.2. Zipformer (Yao et al., 2024), which uses dropout (Srivastava et al., 2014) and stochastic depth (Huang et al., 2016), is used as our speech encoder due to its speed and high performance. Following (Yao et al., 2024), pruned transducer (Kuang et al., 2022), a highly optimized and memory-efficient version of transducer, is employed for comparison. Word-error-rate (WER) and character-error-rate (CER) are employed as ASR metrics for English and Mandarin datasets, respectively. As CR-CTC requires two forward pass during training, we train CR-CTC models with half the batch size and half the number of epochs compared to CTC models, ensuring a fair comparison in terms of training cost. Training configuration in terms of number of GPUs and training epochs are provided in Appendix Section A.2. For CTC and CR-CTC systems, we use prefix search decoding (Graves et al., 2006) with a beam size of 4 for comparisons against other state-of-the-art models, and employ greedy search decoding for ablation studies. Results comparison between these two decoding methods are provided in Appendix Section A.3.
@@ 4.2 Comparison with state-of-the-art models
In this section, we compare our CR-CTC with other state-of-the-art models. For LibriSpeech and GigaSpeech datasets, we also use CR-CTC as an auxiliary loss in CTC/AED and pruned transducer systems for joint training (denoted as CR-CTC/AED and pruned transducer w/ CR-CTC), to further validate the representation learning capability of CR-CTC. For the larger GigaSpeech dataset, we additionally use a even larger scale of Zipformer (Zipformer-XL). Model configuration of different scales of Zipformer are provided in Table 14. For Aishell-1 dataset, which is considerably smaller, we conduct experiments on Zipformer-S and Zipformer-M to ensure comparable parameter counts with other models reported in the literature.
@@
Table 3: WER(%) performance of our method on GigaSpeeech dataset compared to the best results reported in the literature without using an external language model.
@@ 4.3 Ablation studies
Self-distillation. As presented in Table 4, both the lack of increased time masking and the absence of different augmented views lead to WER degradation, indicating the effectiveness of enhancing the input diversity between sub-models during self-distillation. Replacing $D_{\mathrm{KL}}$ with hard label-based cross-entropy (CE) function in $\mathcal{L}_{\mathrm{CR}}$ (Equation 4) results in a WER degradation of 0.02%/0.22% on test-clean/test-other. This suggests the advantage of using $D_{\mathrm{KL}}$ which enables a finer-grained self-distillation as it distills over the full CTC lattice, whereas the hard label CE-based method only distills the best alignment. When removing the sg operation in $\mathcal{L}_{\mathrm{CR}}$, the WER increase by 0.12%/0.35%, which implies that the model might have a tendency towards a degenerated solution (Chen & He, 2021) that is insensitive to the pattern of input masking and model dropout.
@@
Compared to using auxiliary head for jointly training. The straightforward approach to improve the CTC performance is using an auxiliary head of AED (Chan et al., 2015; Watanabe et al., 2017) or pruned transducer (Kuang et al., 2022) for jointly training. As reported in Table 7, CR-CTC significantly outperforms these two methods with less model parameters, suggesting the advantage of our method.
@@ A.1 Smooth-regularized CTC
Smooth-regularized CTC (SR-CTC) discourages peaky distributions by adding an smooth regularization loss (denoted as $\mathcal{L}_{\mathrm{SR}}$) to regular CTC model. Specifically, we first apply a smooth kernel $K$ of size 3 to the model prediction $\mathbf{z}$, smoothing it along the time dimension: $\mathbf{z}^{(s)}=\mathrm{\emph{smooth}}(\mathbf{z},K)$  [^3] The smoothing operation is done by using a 1-D depth-wise convolution layer.. Then we minimize the $D_{\mathrm{KL}}$ between $\mathbf{z}$ and $\mathbf{z}^{(s)}$, similar to the consistency loss in CR-CTC (Equation 4):
@@
$\mathcal{L}^{\prime}=\mathcal{L}_{\mathrm{CTC}}(\mathbf{x},\mathbf{y})+\beta\mathcal{L}_{\mathrm{SR}}(\mathbf{z},\mathbf{z}^{(s)}),$ (6)
@@
We validate its effectiveness in Section 4.3. Table 6 presents the experimental results.
@@ A.3 Results of different decoding methods
Results comparison between greedy search decoding and prefix search decoding for CTC and CR-CTC on LibriSpeech, Aishell-1 and GigaSpeech datasets are presented in Table 11, Table 12, and Table 13, respectively.

### v3

@@ Abstract
Connectionist Temporal Classification (CTC) is a widely used method for automatic speech recognition (ASR), renowned for its simplicity and computational efficiency. However, it often falls short in recognition performance. In this work, we propose the Consistency-Regularized CTC (CR-CTC), which enforces consistency between two CTC distributions obtained from different augmented views of the input speech mel-spectrogram. We provide in-depth insights into its essential behaviors from three perspectives: 1) it conducts self-distillation between random pairs of sub-models that process different augmented views; 2) it learns contextual representation through masked prediction for positions within time-masked regions, especially when we increase the amount of time masking; 3) it suppresses the extremely peaky CTC distributions, thereby reducing overfitting and improving the generalization ability. Extensive experiments on LibriSpeech, Aishell-1, and GigaSpeech datasets demonstrate the effectiveness of our CR-CTC. It significantly improves the CTC performance, achieving state-of-the-art results comparable to those attained by transducer or systems combining CTC and attention-based encoder-decoder (CTC/AED). We release our code at https://github.com/k2-fsa/icefall.
@@
Peaky CTC distributions. CTC models are known for predicting extremely peaky distributions (Graves et al., 2006; Sak et al., 2015), which can be harmful in certain scenarios, such as forced alignment (Huang et al., 2024) and knowledge distillation (Ding et al., 2020). These peaky distributions lead to inaccurate alignments as the model assigns excessive blanks to non-silence frames. To address this, label priors are employed to suppress the peaky distributions, thereby improving the accuracy of forced alignment (Huang et al., 2024). As position mismatches of CTC spikes can hinder knowledge distillation performance, some approaches propose to encourage consistent alignments between the teacher and student (Ding et al., 2020) or to utilize sequence-level distillation (Takashima et al., 2019). Unlike previous works, we demonstrate that peak suppression in CR-CTC can improve the generalization ability of the CTC models.
@@
Consistency regularization. The technique of consistency regularization has demonstrated effectiveness in learning generalizable image representations across various learning paradigms, including self-supervised (Chen et al., 2020; Grill et al., 2020; He et al., 2020; Chen & He, 2021), semi-supervised (Sajjadi et al., 2016; Laine & Aila, 2016; Sohn et al., 2020), and supervised (Wu et al., 2021; Touvron et al., 2023; Heo et al., 2023) learning tasks. Self-supervised methods, such as SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), MoCo (He et al., 2020) and SimSiam (Chen & He, 2021), aim to align hidden representations of unlabeled image data from different model branches or different augmented views. They address the training issue of feature collapsing into a constant vector (Chen & He, 2021) through contrastive learning (Chen et al., 2020; He et al., 2020), momentum encoder (Grill et al., 2020; He et al., 2020), and stop-gradient operation (Chen & He, 2021). In semi-supervised learning, a prominent example leveraging consistency regularization is FixMatch (Sohn et al., 2020). It generates pseudo-labels based on high-confidence predictions from weakly augmented images, then trains the model to predict these pseudo-labels using the strongly augmented versions of the same images. Additionally, in supervised learning, methods such as R-Drop (Wu et al., 2021) and cosub (Touvron et al., 2023) encourage consistency between predictions of randomly sampled sub-models using drop-based techniques.
When employing consistency regularization as unsupervised objective to train transformer encoders on unlabeled speech data, a new training issue arises in the form of the shortcut learning problem (Geirhos et al., 2020), which is tackled using reconstruction loss in Speech SimCLR (Jiang et al., 2020) and temporal augmentation in C-Siam (Khorram et al., 2022). Some studies explore leveraging consistency regularization to enhance model robustness during predicting the pseudo-labels of untranscribed data, which are generated based on different augmentations (Masumura et al., 2020; Weninger et al., 2020; Chen et al., 2021b; Higuchi et al., 2021; Sapru, 2022) or through speech chain reconstruction (Qi et al., 2022). In contrast to these self/semi-supervised ASR works, our work focuses on a fully supervised setting, where we introduce consistency loss as a regularization term to improve performance of CTC model trained on labeled data. As the consistency regularization is enforced on CTC distributions, which are stably supervised by the main CTC loss, it inherently avoids the training issues associated with the unsupervised objectives as observed in Speech SimCLR (Jiang et al., 2020) and C-Siam (Khorram et al., 2022).
The idea of R-Drop (Wu et al., 2021) has also been extended to supervised ASR (Gao et al., 2022; Yoon et al., 2024). For example, to improve the CTC/AED system,  (Gao et al., 2022) specially designs the spatial-temporal dropout to construct the sub-models, with consistency regularization enforced exclusively on the CTC spike frames. Cons-KD (Yoon et al., 2024) integrates consistency regularization into a knowledge distillation system, enabling the student model to be more robust to inconsistency induced by dropout. In this work, we focus on improving the performance of pure CTC systems and are the first to enable CTC models to match the performance of transducer and CTC/AED systems by a simple yet effective approach. Moreover, we introduce peak suppression as a novel explanatory perspective, demonstrating for the first time that it can mitigate overfitting and enhance the generalization ability of CTC models.
@@ 3.1 Preliminary: Connectionist Temporal Classification
$\vspace{-0.4em}\mathcal{L}_{\mathrm{CTC}}(\mathbf{x},\mathbf{y})=-\log\sum_{\bm{\pi}\in\mathcal{B}^{-1}(\mathbf{y})}p(\bm{\pi}|\mathbf{x}).$ (1)
@@ 3.2 Our approach: Consistency-regularized CTC
Figure 1 illustrates the overall architecture of our proposed CR-CTC. It takes as input two different augmented views, $\mathbf{x}^{(a)}$ and $\mathbf{x}^{(b)}$, both derived from the input speech mel-spectrogram $\mathbf{x}$. The two input views are then passed through a shared speech encoder $f$, which estimates the per-frame distributions: $\mathbf{z}^{(a)}=f(\mathbf{x}^{(a)})$, $\mathbf{z}^{(b)}=f(\mathbf{x}^{(b)})$. In addition to computing the CTC losses on both branches: $\mathcal{L}_{\mathrm{CTC}}(\mathbf{z}^{(a)},\mathbf{y})$ and $\mathcal{L}_{\mathrm{CTC}}(\mathbf{z}^{(b)},\mathbf{y})$, we introduce an auxiliary loss (defined in Equation 4) to enforce consistency between $\mathbf{z}^{(a)}$ and $\mathbf{z}^{(b)}$: $\mathcal{L}_{\mathrm{CR}}(\mathbf{z}^{(a)},\mathbf{z}^{(b)})$. The overall loss of the whole model is formulated as:
$\mathcal{L}=\frac{1}{2}(\mathcal{L}_{\mathrm{CTC}}(\mathbf{z}^{(a)},\mathbf{y})+\mathcal{L}_{\mathrm{CTC}}(\mathbf{z}^{(b)},\mathbf{y}))+\alpha\mathcal{L}_{\mathrm{CR}}(\mathbf{z}^{(a)},\mathbf{z}^{(b)}),$ (3)
@@ 4.1 Experimental setup
Implementation details. Our experiments are performed using the icefall framework  [^2] https://github.com/k2-fsa/icefall, with Lhotse toolkit (Żelasko et al., 2021) for data preparation. For regular ASR recipes in icefall, default parameter settings of SpecAugment (Park et al., 2019) include a time warping factor of 80, 2 frequency masking regions with a maximum width of 27, and 10 time masking regions with a maximum width of 100, along with a maximum masking fraction of 15% specifically for time masking  [^3] See the SpecAugment implementation in Lhotse for more details: https://github.com/lhotse-speech/lhotse/blob/master/lhotse/dataset/signal_transforms.py. In our CR-CTC systems, we utilize larger amount of time masking through increasing both the number of time masking regions and the maximum masking fraction by a factor of 2.5. Speed perturbation (Ko et al., 2015) with factors 0.9, 1.0 and 1.1 is applied to LibriSpeech (Panayotov et al., 2015) and Aishell-1 (Bu et al., 2017) datasets. The input features are 80-dimensional mel-spectrograms extracted using 25-ms windows with a 10-ms shift. For LibriSpeech and GigaSpeech datasets, we employ 500-class Byte Pair Encoding (BPE) (Sennrich et al., 2016) word pieces as modeling units, while for Aishell-1 dataset, we use 4336-class characters. By default, we set $\alpha$ in Equation 3 to 0.2. Zipformer (Yao et al., 2024), which uses dropout (Srivastava et al., 2014) and stochastic depth (Huang et al., 2016), is used as our speech encoder due to its speed and high performance. It takes input features at frame rate of 100Hz, processes the sequence through 6 stacks with frame rates of 50Hz, 25Hz, 12.5Hz, 6.25Hz, 12.5Hz, and 25Hz, and finally produces the encoder output at frame rate of 25Hz. Following (Yao et al., 2024), pruned transducer (Kuang et al., 2022), a highly optimized and memory-efficient version of transducer, is employed for comparison. Word-error-rate (WER) and character-error-rate (CER) are employed as ASR metrics for English and Mandarin datasets, respectively. As CR-CTC requires two forward pass during training, we train CR-CTC models with half the batch size and half the number of epochs compared to CTC models, ensuring a fair comparison in terms of training cost. Training configuration in terms of number of GPUs and training epochs are provided in Appendix Section A.2. For CTC and CR-CTC systems, we use prefix search decoding (Graves et al., 2006) with a beam size of 4 for comparisons against other state-of-the-art models, and employ greedy search decoding for ablation studies. Results comparison between these two decoding methods are provided in Appendix Section A.3. For pruned transducer models, we use beam search decoding with beam size of 4 (Kang et al., 2023). For CTC/AED systems, we use joint decoding that combines CTC scores and AED scores (Watanabe et al., 2017).
@@ 4.2 Comparison with state-of-the-art models
In this section, we compare our CR-CTC with other state-of-the-art models. For LibriSpeech and GigaSpeech datasets, we also use CR-CTC as an auxiliary loss in CTC/AED and pruned transducer systems for joint training (denoted as CR-CTC/AED and pruned transducer w/ CR-CTC), to further validate the representation learning capability of CR-CTC. Note that for the models that combine CR-CTC and pruned transducer, we only utilize the transducer head for decoding, without incorporating the CTC scores. For the larger GigaSpeech dataset, we additionally use a even larger scale of Zipformer (Zipformer-XL). Model configuration of different scales of Zipformer are provided in Table 15. For Aishell-1 dataset, which is considerably smaller, we conduct experiments on Zipformer-S and Zipformer-M to ensure comparable parameter counts with other models reported in the literature.
@@
Table 3: WER(%) performance of our method on GigaSpeech dataset compared to the best results reported in the literature without using an external language model.
@@ 4.3 Ablation studies
Self-distillation. One self-distillation method in self-supervised learning is to construct a teacher model by tracking the model weights using exponential moving average (EMA) (Grill et al., 2020; He et al., 2020; Baevski et al., 2023). For comparison, we include this approach, referred to as EMA-distilled CTC, which incorporates an auxiliary loss to learn from the CTC distribution of the EMA teacher model. Its details are provided in Appendix Section A.6. As presented in Table 4, CR-CTC significantly outperforms EMA-distilled CTC, demonstrating its superiority in self-distillation. For CR-CTC, both the lack of increased time masking and the absence of different augmented views lead to WER degradation, indicating the effectiveness of enhancing the input diversity between sub-models during self-distillation. Replacing $D_{\mathrm{KL}}$ with hard label-based cross-entropy (CE) function in $\mathcal{L}_{\mathrm{CR}}$ (Equation 4) results in a WER degradation of 0.02%/0.22% on test-clean/test-other. This suggests the advantage of using $D_{\mathrm{KL}}$ which enables a finer-grained self-distillation as it distills over the full CTC lattice, whereas the hard label CE-based method only distills the best alignment. When removing the sg operation in $\mathcal{L}_{\mathrm{CR}}$, the WER increase by 0.12%/0.35%, which implies that the model might have a tendency towards a degenerated solution (Chen & He, 2021) that is insensitive to the pattern of input masking and model dropout.
@@
Compared to using auxiliary head for jointly training. The straightforward approach to improve the CTC performance is using an auxiliary head of AED (Chan et al., 2015; Hentschel et al., 2024) or pruned transducer (Kuang et al., 2022) for jointly training, while retaining only the CTC head for inference. As reported in Table 7, CR-CTC significantly outperforms these two methods with less model parameters, suggesting the advantage of our method.
@@ A.1 Smooth-regularized CTC
Smooth-regularized CTC (SR-CTC) discourages peaky distributions by adding an smooth regularization loss (denoted as $\mathcal{L}_{\mathrm{SR}}$) to regular CTC model. Specifically, we first apply a smooth kernel $K$ of size 3 to the model prediction $\mathbf{z}$, smoothing it along the time dimension: $\mathbf{z}^{(s)}=\mathrm{\emph{smooth}}(\mathbf{z},K)$. The smoothing operation is done by using a 1-D depth-wise convolution layer. Then we minimize the $D_{\mathrm{KL}}$ between $\mathbf{z}$ and $\mathbf{z}^{(s)}$, similar to the consistency loss in CR-CTC (Equation 4):
@@
$\mathcal{L}^{\prime}=\mathcal{L}_{\mathrm{CTC}}(\mathbf{z},\mathbf{y})+\beta\mathcal{L}_{\mathrm{SR}}(\mathbf{z},\mathbf{z}^{(s)}),$ (6)
@@
We validate its effectiveness in Section 4.3. Table 6 presents the experimental result.
@@ A.3 Results of different decoding methods
Results comparison between greedy search decoding and prefix search decoding for CTC and CR-CTC on LibriSpeech, Aishell-1 and GigaSpeech datasets are presented in Table 11, Table 12, and Table 13, respectively. In addition, Table 14 presents the results of different beam sizes for prefix search decoding on LibriSpeech dataset with Zipformer-M encoder.
@@
Table 14: WER (%) results of different beam sizes for prefix search decoding on LibriSpeech dataset using Zipformer-M encoder.
Method Beam size test-clean test-other
CTC 1 2.52 6.02
2 2.52 6.02
4 2.52 6.02
8 2.52 6.02
CR-CTC 1 2.1 4.61
2 2.1 4.61
4 2.1 4.61
8 2.1 4.61
@@ A.5 Ablation Studies on Hyper-parameter Tuning
A.6 EMA-distilled CTC
In EMA-distilled CTC, the teacher model $f^{(e)}$ is dynamically constructed for self-distillation. Its weights $\theta^{(e)}$ are updated using the exponential moving average of the current model’s weights $\theta$: $\theta^{(e)}\leftarrow\tau\theta^{(e)}+(1-\tau)\theta$, where $\tau=\min(0.9999,1-10/\max(20,\mathrm{step}))$. The teacher model $f^{(e)}$ processes the unmasked input $\mathbf{x}^{(e)}$, and produces the CTC distribution $\mathbf{z}^{(e)}=f^{(e)}(\mathbf{x}^{(e)})$ which serves as distillation target for the current model $f$. Similar to $\mathcal{L}_{\mathrm{CR}}$ in CR-CTC (Equation 4), the distillation loss $\mathcal{L}_{\mathrm{EMA}}$ is defined as:
$\mathcal{L}_{\mathrm{EMA}}(\mathbf{z},\mathbf{z}^{(e)})=\sum_{t=1}^{T}D_{\mathrm{KL}}(\mathrm{\emph{sg}}(z^{(e)}_{t})\|z_{t}).$ (7)
The overall loss of EMA-distilled CTC is formulated as:
′′$\mathcal{L}^{\prime\prime}=\mathcal{L}_{\mathrm{CTC}}(\mathbf{z},\mathbf{y})+\gamma\mathcal{L}_{\mathrm{EMA}}(\mathbf{z},\mathbf{z}^{(e)}),$ (8)
where $\gamma$ is a hyper-parameter. In this work, we use $\gamma=0.2$. Table 4 presents the experimental result.
A.7 Results using Conformer encoder
To validate the effectiveness and generalization ability of our proposed CR-CTC, we conduct experiments on LibriSpeech dataset using a Conformer (Gulati et al., 2020) encoder, comparing different methods including CTC (Graves et al., 2006), CTC/AED (Watanabe et al., 2017), pruned transducer (Kuang et al., 2022) and CR-CTC. Specifically, we use a 12-layer Conformer, with an embedding dimension of 512, a convolution kernel size of 31, and a feedforward hidden dimension of 2048. For the CTC/AED model, the AED decoder is modeled by a 6-layer Transformer, where each layer has an attention dimension of 512 and a feedforward hidden dimension of 2048. The vanilla CTC model is trained for 100 epochs, while the other three models are trained for 50 epochs. Table 17 presents the experimental results. CR-CTC substantially outperforms the vanilla CTC and achieves marginally better results compared to the pruned transducer and CTC/AED.
Table 17: WER(%) performance of difference methods on LibriSpeech dataset using a 12-layer Conformer encoder.
Method Params (M) WER (%)
test-clean test-other
CTC 77.4 2.92 7.15
CTC/AED 103.1 2.5 5.94
Pruned transducer 78.6 2.49 5.87
CR-CTC (ours) 77.4 2.43 5.78